# Latent Beam Diffusion Models for Decoding Image Sequences

### v2

@@ 5 Results and Discussion
In this section, we describe the performance of BeamDiffusion based on both human and automatic evaluations. We begin by analyzing the human evaluation, proceed to discussing the automatic evaluations, then focus on specifics of our approach such as hyperparameters and non-linear decoding characteristics, and finally provide a qualitative analysis. Additional statistical information about the results can be found in Appendix B.1.
@@ 5.1 General Results
Table 1 compares all methods across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain). BeamDiffusion consistently outperforms other methods. In the context of Recipes, BeamDiffusion is selected 50.0% of the time by humans and 37.6% by the LLM, with Nucleus and Greedy following. In out-of-domain areas, it is chosen once more 50.0% of the time by humans and 30.0% by the LLM-judge for the DIY domain, and 58.3% for humans and 34.9% by the LLM for the VIST dataset. Among sequence generation methods, GILL consistently underperforms across all domains, receiving at most 3.3% of human votes and 16.1% from the LLM in VIST. StackedDiffusion performs better than both Greedy and GILL, but still lags behind BeamDiffusion, with 20.0% of human selections and 9.8% of LLM-judge selections in the Recipes domain.
Overall, BeamDiffusion excels in Recipes and VIST, however performance in DIY remains challenging, with smaller gaps between methods due to the domain’s inherent complexity.
Table 1: Frequency of method selection by human annotators(H) and Gemini(G).
Recipes DIY VIST
Method H (%) G (%) H (%) G (%) H (%) G (%)
GILL [18] 0.0 2.0 0.0 20.0 3.3 16.1
StackedDiffusion [26] 20.0 9.8 12.5 15.6 0.00 3.2
Image Sequence Decoding
Greedy/CoSeDlen=1 [33] 0 27.6 0 23.3 0 26.3
Nucleus Sampling [11] 30.0 23.0 37.5 27.7 8.3 26.6
BeamDiffusion 50.0 37.6 50.0 30.0 58.3 34.9
@@ 5.2 Automatic Metrics
Table 2 shows that BeamDiffusion achieves the highest CLIP* and DINO* scores (0.40 and and 0.30), demonstrating a strong balance between image consistency and semantic alignment. BeamDiffusion excels in maintaining a good compromise between text grounding and image similarity achieving the best result in image consistency (CLIP-I of 0.83 and DINO-I of 0.62). Similarly, for CLIP-T, BeamDiffusion achieves the highest score for image-text alignment (CLIP-T of 0.48). However, in this case, the highest result is shared, as Greedy and Nucleus obtain the same score. This outcome is expected because Greedy and Nucleus utilize CLIP-T to select their outputs by aligning the generated sequence with the text. In terms of sequence generation methods, StackedDiffusion performs better than GILL in all metrics but still lags behind BeamDiffusion and the other decoding methods.
Table 2: Comparison of CLIP-I, DINO-I and CLIP-T across different methods
Method CLIP-I DINO-I CLIP-T CLIP* DINO*
GILL 0.55 0.49 0.43 0.23 0.21
StackedDiffusion 0.76 0.46 0.47 0.36 0.22
Image Sequence Decoding
Greedy 0.81 0.54 0.48 0.39 0.26
Nucleus 0.77 0.53 0.48 0.37 0.25
BeamDiffusion 0.83 0.62 0.48 0.40 0.30
5.3 Best Beam Search Configuration
@@
We first perform human annotations on a set of sequences to identify the optimal beam search configuration, see Appendix B.1 for details. To determine the best configuration, annotations were used to discard non-optimal configurations. The results in Table 4 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%. Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed a strong performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration. The configurations with beam width 6 performed the worst in all the settings tested.
To assess the reliability of these findings, we used Gemini for the same evaluation on six tasks, achieving an 85% Fleiss’ Kappa agreement [6], indicating near-perfect alignment. The results from the extended Gemini evaluation (Table 4) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained the best, now achieving a 48.57% selection rate, a significant improvement over the 33.33% from the initial six-task evaluation. Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative. However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
@@
Steps Back Beam Width
@@
Table 4: Influence of different hyperparameter values according to Gemini annotations.
Steps Back Beam Width
@@
5.4 Non-Linear Sequence Denoising
To analyze how BeamDiffusion explores the latent space, we measured the number of times that a sequence step provides latents for the generation of subsequent steps of the sequence. In this experiment we consider sequences of four steps, (Figure 2). Appendix C.2 provides an analysis at the level of the denoising iterations. Latents from step 1 were chosen 69.9% of
Figure 2: Non-linear steps selection, according to selection opportunities.
the times they had a chance of being selected, highlighting the role of step 1 in shaping the sequence. Selection rates decrease for later steps, with Step 3 at 46.8% and Step 2 at 33.3%. Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate. Thus, the percentages reflect how often each step is selected given the different possibilities for their inclusion, rather than a simple additive total. This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate sequence steps.
(a) Recipes domain task.
(b) DIY task.
(c) VIST story.
Figure 3: Example tasks from three different domains: Recipes ("Jamaican Callaloo With Shrimp"), DIY:("How to Make a Door"), and VIST ("July 4th Barbecue").
5.5 Qualitative Analysis
To highlight how different decoding methods and models impact the semantic and visual consistency of generated outputs, we present several examples in Figures 3(a)3(c) and Appendix F.2. We compare three decoding methods (Greedy, Nucleus, and Beam) with two models (GILL and StackedDiffusion).
In Figure 3(a), BeamDiffusion consistently maintains the same pan across all steps, a feature not always seen in Greedy or Nucleus methods. While GILL struggles to follow the step text and results in more inconsistencies, Beam performs better in preserving visual coherence compared to Greedy and Nucleus, though it still faces some challenges. For DIY tasks,as shown in Figure 3(b), BeamDiffusion can preserve basic elements such as the color of the plywood and a basic adherence to the text. However, overall, all methods face challenges in capturing precise details. In the VIST dataset, Figure 3(c), Greedy and Nucleus alternate between black-and-white and color images, while BeamDiffusion ensures color consistency throughout the sequence. Furthermore, background elements, such as bushes, are maintained in the first three steps, and the individuals in the first and third steps closely resemble each other. Overall, BeamDiffusion excels in maintaining both semantic and visual consistency, even in complex out-of-domain tasks, In contrast, methods such as StackedDiffusion and GILL encounter challenges in keeping a balance between text alignment and consistency across images.
6 Limitations
The theoretical complexity of BeamDiffusion is linear with the number of beams that we configure the model to explore. The drawback of this strategy is that exploring the space of hypothesis increases the computational complexity of the process. However, this provides a solid ground to explore exciting new ideas where stronger constraints can be enforced in the explored beams.
7 Conclusions
In this paper, we proposed a novel method to explore the latent diffusion space with a beam search strategy, to generate image sequences. BeamDiffusion is a theoretically sound framework, drawing on the strengths of text-based decoding techniques [40], with clear experimental gains over alternative methods, as we demonstrated in generating image sequences with enhanced coherence and consistency.
Through both human and automatic evaluation, we show that beam search performs effectively in generating high-quality and coherent image sequences. The insights drawn from these evaluations reveal that beam search aligns closely with textual instructions while ensuring visual continuity. Most methods can align images with text, however, we demonstrated that sequence decoding methods perform naturally better at aligning their output with the input sequence, and BeamDiffusion in particular achieves an optimal balance between the sequence of text instructions and the visual coherence of the generated images. Furthermore, our results underscore the significance of the non-linear selection process in beam search, highlighting how prioritizing certain steps and latents can optimize the quality of the generation process.
This work shows that the latent diffusion space can be steered, offering more reliable and factual outputs. As future work, we plan to explore constraint decoding method that assess the likelihood of each denoise latent with respect to the target prompt. Solving this challenge, will improve the correlation between all elements in a sequence of scenes.
Appendix A Implementation Details
Base Model.
Our method is implemented on top of Stable Diffusion v2.1 [35], a latent diffusion model for high-quality text-to-image synthesis. We build upon the open-source implementation provided by Stability AI, extending it to support beam-based generation, latent trajectory tracking, and contrastive scoring mechanisms as described in the main text.
Latent Trajectory Reuse.
Classifier Training.
To train the contextual classifier, we employ a sequence-aware training approach that allows the model to capture dependencies between steps in the generation process. The classifier $\varphi$ takes as input a candidate image $I_{l}$, its textual description $s_{l}$, and the previous sequence images and the respective step text ($B_{l-1}^{(b)}$). Then the classifier scores $I_{j}$ based on how well it fits within the evolving trajectory. During training, we use labels $l_{d,l}$ indicating whether $I_{l}$ is the correct continuation in the sequence for task $d$ and step $l$. The model is trained to minimize the cross-entropy loss:
Hardware.
All experiments were ran on NVIDIA A100 GPUs with 40GB of memory.
Appendix B Experiments Details
B.1 Metrics and Annotations

### v3

@@ 5 Results and Discussion
In this section, we evaluate the effectiveness of BeamDiffusion through three complementary approaches: human evaluations, LLM-as-a-judge, and automatic metrics. We begin by comparing decoding strategies, selecting the method that achieves the highest human preference scores for subsequent experiments. We then highlight key performance aspects of BeamDiffusion, including visual and semantic consistency, and conclude with a qualitative analysis. Additional details on the annotation setup are provided in Appendix E.
5.1 Comparison of Decoding Methods
To evaluate the effectiveness of different decoding strategies, we compare our proposed method, BeamDiffusion, against Greedy/CoSeDlen=1 [37] and Nucleus Sampling [13] using the same sequences of prompts with two types of diffusion backbones.
Table 1: Comparison of human (H) and LLM-as-a-judge (L) evaluations of decoding strategies across two backbones.
FLUX.1-dev SD2.1
Method H (%) L (%) H (%) L (%)
Greedy/CoSeDlen=1 [37] 16.35 10.00 11.11 10.00
Nucleus Sampling [13] 24.04 16.67 33.33 26.67
BeamDiffusion 59.61 73.33 55.56 63.33
As shown in Table 1, BeamDiffusion consistently achieves the highest preference scores according to both human annotators (H) and LLM-as-a-judge (L). In particular, BeamDiffusion substantially outperforms Greedy decoding and Nucleus Sampling, indicating that beam-based decoding produces sequences that are not only more coherent but also better aligned with human and automatic judgments. While Nucleus Sampling improves over Greedy decoding by encouraging diversity, its performance remains notably below that of BeamDiffusion in both evaluation settings. The agreement between H and L further strengthens the evidence for the robustness of our decoding strategy. Given these results, we adopt BeamDiffusion as our default decoding method in subsequent experiments (see Appendix C.4 for qualitative and automatic evaluation results and Appendix B for hyperparameter configurations).
Table 1 also compares two diffusion architectures: Stable Diffusion (SD2.1) and Diffusion in Transformers (FLUX-1-dev). BeamDuffusion exhibits consistent performance with both architecture, with Diffusion in Transformers outperforming SD2.1. Hence, in the following experiments we will use the Diffusion in Transformers architecture.
Table 2: Frequency of method selection by human annotators (H) and the LLM-as-a-judge (L).
Recipes VIST
Method H (%) L (%) H (%) L (%)
GILL [20] 0.00 0.00 2.74 0.00
StackedDiffusion [29] 0.00 3.33 0.00 0.00
StoryDiffusion [55] 22.58 20.00 36.99 41.38
1P1S [24] 12.90 3.33 19.18 10.34
BeamDiffusion 64.52 73.33 41.10 48.28
@@ 5.2 General Results
Table 2 compares the sequence generation methods on both the Recipes and VIST datasets. BeamDiffusion dominates in both domains, leading Recipes with 64.52% (H) and 73.33% (L), and VIST with 41.10% (H) and 48.28% (L). StoryDiffusion consistently ranks second, with 22.58% (H) and 20.00% (L) in Recipes, and 36.99% (H) and 41.38% (L) in VIST. In contrast, 1P1S receives moderate preference (12.90% H in Recipes, 19.18% H in VIST), while GILL and StackedDiffusion record the lowest selection rates.
These results show that BeamDiffusion achieves the strongest performance across settings, with especially large margins in Recipes where BeamDiffusion surpasses the next-best method by over 40% for human preference. The consistently low selection rates of GILL and StackedDiffusion highlight that these methods are rarely preferred by either humans or LLM-as-a-judge, in contrast to the stronger performance of BeamDiffusion and StoryDiffusion. Overall, the findings highlight that effective search and selection strategies are central to producing coherent and engaging narratives, enabling models to maintain quality even in complex domains such as VIST.
Table 3: Comparison of different methods across embedding-based sequence quality metrics. CLIP-I, DINO-I, CLIP-T, CLIP*, and DINO* are reported as win percentages, while Goal Faithfulness, Step Faithfulness, and Cross-Image Consistency are reported as absolute scores.
[…]
GILL [20] 0.00 0.00 0.00 0.00 0.00 0.67 0.52 0.54
StackedDiffusion [29] 25.00 25.00 5.00 25.00 25.00 0.60 0.67 0.57
StoryDiffusion [55] 20.00 25.00 35.00 25.00 20.00 0.78 0.87 0.61
1P1S [24] 25.00 30.00 10.00 10.00 25.00 0.75 0.83 0.57
BeamDiffusion 30.00 20.00 50.00 40.00 30.00 0.82 0.92 0.61
@@ 5.3 Automatic Metrics
Table 3 compares all methods using metrics that measure sequence quality, including CLIP-I, DINO-I, CLIP-T, CLIP*, DINO*, and Goal Faithfulness [29], Step Faithfulness [29], and Cross-Image Consistency [29].
BeamDiffusion achieves the highest combined scores (CLIP*: 40.00%, DINO*: 30.00%), reflecting strong performance in both visual and semantic consistency. This is further confirmed by faithfulness and consistency metrics, where BeamDiffusion leads in Goal Faithfulness (0.82), Step Faithfulness (0.92), and Cross-Image Consistency (0.61).
StoryDiffusion performs well on CLIP-T (35.00%) and achieves high scores on the alignment metrics Goal Faithfulness: 0.78, Step Faithfulness: 0.87, Cross-Image Consistency: 0.61, showing strong semantic alignment and visual coherence despite lower combined metrics. 1P1S achieves the highest DINO-I (30.00%) but lower CLIP-T (10.00%) and moderate alignment scores, indicating strong visual consistency but weaker semantic alignment. StackedDiffusion maintains balanced performance across CLIP-I, DINO-I, and CLIP* (around 25.00%) with moderate alignment scores, while GILL consistently underperforms across both automatic and alignment metrics.
Overall, the relationships between CLIP-I/DINO-I and Cross-Image Consistency, and between CLIP-T and Goal/Step Faithfulness, show that the combined metrics CLIP* and DINO* reliably reflect both semantic and visual consistency. Together, these results demonstrate that BeamDiffusion is the most effective method for generating coherent and semantically faithful sequences of images.
5.4 Non-Linear Sequence Denoising
To analyze how BeamDiffusion explores the latent space, we measured the number of times that a sequence step provides latents for the generation of subsequent steps of the sequence. In this experiment we consider sequences of four steps.
Latents from step 1 were chosen 69.9% of the times they had a chance of being selected, highlighting the role of step 1 in shaping the sequence. Selection rates decrease for later steps, with Step 3 at 46.8%, Step 2 at 33.3%, and the random seed at 15.6%. Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate. Thus, the percentages reflect how often each step is selected given the different possibilities for their inclusion, rather than a simple additive total. This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate sequence steps. Appendix C.2 provides mode details.
5.5 Effect of Latent History
Figure 3: Visual comparison of sequences generated with cumulative latent access. The first row shows a sequence generated using latents only from step 1, the second row shows a sequence generated using latents from steps 1 through 5, and the bottom row shows the result from BeamDiffusion, which uses latents from only the two preceding steps.
To better understand the role of latent reuse, we compared three settings: using latents only from step 1, using latents from all steps (1–5), and using BeamDiffusion (Figure 3). Sequences generated from step 1 alone lacked detail and consistency, while access to all steps improved quality but came with higher computational cost. In contrast, BeamDiffusion, which relies only on the two most recent steps, achieved results comparable to or better than full access while being substantially more efficient. This efficiency arises because each new latent already encodes information from previous steps (Figure 2(b)), making full-history conditioning unnecessary. A more detailed analysis of intermediate cases, is provided in Appendix C.3.
(a) Recipes domain task.
(b) VIST story.
Figure 4: Example tasks from two different domains: Recipes ("Crema Catalana with salted caramel sauce") and VIST ("Fireworks - 4th July in Washington, DC").
5.6 Qualitative Analysis
To highlight the impact of our approach on semantic and visual consistency, we present two examples, one from the Recipes domain (Figure 4(a)) and one from VIST (Figure 4(b)). In the recipe example, BeamDiffusion avoids major semantic errors and maintains consistent object shapes across steps; for instance, while StoryDiffusion and 1P1S show inconsistencies in the shape of the tins between steps 3 and 4, BeamDiffusion preserves them accurately. In the VIST example, BeamDiffusion maintains both semantic alignment and visual coherence throughout the sequence, accurately depicting two friends and preserving consistent colors and lighting across all steps. In contrast, baselines such as 1P1S introduce a temporal inconsistency, shifting from sunset to night and back to sunset, while StoryDiffusion misinterprets the step text by showing three individuals instead of two friends.
6 Limitations
The theoretical complexity of BeamDiffusion is linear with the number of beams that we configure the model to explore. The drawback of this strategy is that exploring the space of hypothesis increases the computational complexity of the process, discussed in Appendix G. However, this provides a solid ground to explore exciting new ideas where stronger constraints can be enforced in the explored beams.
7 Conclusions
In this paper, we proposed a novel method to explore the latent diffusion space with a beam search strategy, to generate image sequences. BeamDiffusion is a theoretically sound framework, drawing on the strengths decoding techniques [44], with clear experimental gains over alternative methods, as we demonstrated in generating image sequences with enhanced coherence and consistency.
Through both human and automatic evaluation, we show that beam search performs effectively in generating high-quality and coherent image sequences. The insights drawn from these evaluations reveal that beam search aligns closely with textual instructions while ensuring visual continuity. Most methods can align images with text, however, we demonstrated that sequence decoding methods perform naturally better at aligning their output with the input sequence, and BeamDiffusion in particular achieves an optimal balance between the sequence of text instructions and the visual coherence of the generated images. Furthermore, our results underscore the significance of the non-linear selection process in beam search, highlighting how prioritizing certain steps and latents can optimize the quality of the generation process. This work shows that the latent diffusion space can be steered, offering more reliable and factual outputs. As future work, we plan to explore constraint decoding method that assess the likelihood of each denoise latent with respect to the target prompt. Solving this challenge, will improve the correlation between all elements in a sequence of scenes.
Appendix A Contextual Scene Descriptions
In many cases, step descriptions alone may lack context as precise visual prompts. These step descriptions are often interdependent, relying on prior steps for contextual understanding, or may miss visual details, e.g. "Put the ice cream in freezer for 1 hour", or actions that involve multiple tasks, such as "Slice the tomatoes, grate the cheese, and stir the sauce". To address this challenge, we leverage Gemini [38], specifically the Gemini 1.5 Flash model, to refine step descriptions into visually detailed prompts while considering sequence context. Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
$c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (6)
Appendix B Best Beam Search Configuration
@@
We first perform human annotations on a set of sequences to identify the optimal beam search configuration, see Appendix E for details. To determine the best configuration, annotations were used to discard non-optimal configurations. The results in Table 5 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%. Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed a strong performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration. The configurations with beam width 6 performed the worst in all the settings tested.
To verify the reliability of our findings, the LLM-as-a-judge was applied to six tasks, yielding an 85% Fleiss’ Kappa agreement [8], reflecting near-perfect alignment. The extended LLM-as-a-judge evaluation (Table 5) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained best, reaching a 48.57% selection rate compared to 33.33% in the initial six-task evaluation. Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative. However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
@@
Steps
Back
Beam Width
@@
Table 5: Influence of different hyperparameter values according to LLM-as-a-judge annotations.
Steps
Back
Beam Width
@@
Appendix C Impact of Latent Selection on Beam Search Performance
C.1 Impact of Latent Groups on Performance
Following the initial analysis of beam search configurations (Appendix B), where we used latent indexes 0, 1, 2, and 3, we further investigated the impact of using different latent configurations on the generated sequences. While latents 0, 1, 2, and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance. In particular, beam search configurations for these underperforming latents result in very limited selection rates, indicating they are rarely chosen in both human and LLM-as-a-judge annotations. For human annotations across 6 tasks, the only configuration that yielded any selection was the one with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%. All other configurations resulted in 0% selection. In contrast, none of the configurations showed any non-zero selection rate under LLM-as-a-judge annotations, reinforcing the lack of preference for these latents.
Figure 5: Impact of using late-index latents configurations on sequences generation.
Similarly, for LLM-as-a-judge annotations evaluated on the same 6 sequences, the configurations do not yield any notable results, with no selection rates. When the evaluation is extended to 35 sequences, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%. Figure 5 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [37, 3].
C.2 Latent Selection on Image Sequence Generation
C.3 Effect of Cumulative Latent Access
As shown in Figure 7, including latents from only the first step makes step 4 appear very closed up. Limiting the model to latents from the first two steps makes it more difficult to follow the step text while maintaining coherence, which is particularly evident in step 5. Gradually expanding the latent set to include more steps improves sequence quality, especially when moving from two to three steps. However, including latents from all previous steps does not yield further improvements in coherence, suggesting diminishing returns with increasing temporal access.
In contrast, BeamDiffusion uses latents from only the two most recent steps and achieves results that are comparable, or even superior, to the full cumulative case. The reason is that sequence generation naturally forms a latent chain: to generate a new image, the model conditions on a latent from the previous step, which has already absorbed information from earlier generations. The resulting latent is then updated with the new step, effectively carrying forward both old and new information (see Figure 2(b) for an illustration of this chain effect). In this way, the most recent latents act as a compressed summary of the entire history, making access to all earlier latents redundant. By always using only the last two latents, BeamDiffusion maintains a global summary of the sequence while preserving efficiency.
Overall, this experiment highlights that cumulative latent access enhances generation quality up to a point, but full historical memory is unnecessary. Strategies such as BeamDiffusion, which leverage a limited yet informative temporal window, can produce high-quality sequences efficiently. These findings also underscore the importance of the non-linear sequence denoising strategy introduced in Section 5.4, which enables more targeted and effective reuse of past latents.
C.4 Automatic and Qualitative Evaluation of Decoding Strategies
Table 6: Comparison of different decoding methods across embedding-based sequence quality metrics. CLIP-I, DINO-I, CLIP-T, CLIP*, and DINO* are reported as win percentages, while Goal Faithfulness, Step Faithfulness, and Cross-Image Consistency are reported as absolute scores.
Method CLIP-I DINO-I CLIP-T CLIP* DINO*
[…]
Greedy/CoSeDlen=1 30.00 30.00 35.00 25.00 30.00 0.76 0.90 0.62
Nucleus Sampling 25.00 40.00 40.00 35.00 35.00 0.72 0.80 0.60
BeamDiffusion 45.00 30.00 25.00 40.00 35.00 0.82 0.92 0.61
Figure 8: Illustrative comparison of decoding approaches. BeamDiffusion prevents cascading inconsistencies across steps while preserving semantic and visual consistency.
Automatic evaluation.
Table 6 reports automatic comparisons across similarity metrics and faithfulness/consistency measures. The combined scores CLIP* and DINO* offer a more holistic signal than their individual components. Here, BeamDiffusion achieves the best CLIP* score (40.00) and matches the highest DINO* (30.00), indicating strong performance across both text–image and image–image similarity. By contrast, Greedy/CoSeD lags significantly on CLIP* (25.00) and only reaches 30.00 on DINO*, while Nucleus Sampling improves slightly on CLIP* (35.00) and DINO* (35.00) but at the expense of weaker instruction alignment and coherence.
BeamDiffusion also obtains the strongest scores on goal faithfulness (0.82) and step faithfulness (0.92), showing that its generations remain closely aligned with both the overall task and individual step instructions. Greedy/CoSeD achieves reasonably high step faithfulness (0.90) and the best cross-image consistency (0.62), but falls behind in goal alignment. Nucleus Sampling, in contrast, is weaker across all three measures (0.72, 0.80, 0.60). Taken together, the similarity metrics and the faithfulness/consistency measures highlight that BeamDiffusion avoids the sharp trade-offs of the other decoding methods, achieving a favorable balance between multimodal similarity, instruction following, and visual coherence.
Qualitative evaluation.
Figures 8 and 9 illustrate characteristic error modes of baseline strategies. Nucleus Sampling often suffers from blur propagation, where low-frequency artifacts introduced in one step persist across subsequent generations, degrading the sequence. Greedy decoding avoids blur but instead produces visual inconsistencies, with objects abruptly shifting style or geometry across steps (e.g., the changing appearance of the pan between step 1 and step 3, in Figure 8).
In contrast, BeamDiffusion avoids both pitfalls. By dynamically selecting informative latent trajectories, it prevents error accumulation while preserving smooth semantic and visual transitions. The generated sequences remain sharper, semantically faithful, and more coherent over time. This demonstrates that decoding is not merely a technical detail but a decisive factor in determining the fidelity and consistency of long-horizon generations.
Figure 9: Illustrative comparison of decoding approaches. BeamDiffusion prevents cascading inconsistencies across steps while preserving semantic and visual consistency.
Summary.
Overall, both automatic and qualitative results show that BeamDiffusion delivers a principled improvement over standard strategies. By explicitly managing latent trajectories, it achieves robustness against error propagation and inconsistency while preserving semantic faithfulness and cross-step coherence.
Appendix D Minimizing Decoding Risk
Greedy Decoding and Nucleus Sampling.
BeamDiffusion.
Random Seeds for mid-sequence Generation.
Appendix E Metrics and Annotations